solve the system of equations by the substitution method to get an ordered pair
y=2x+7
y=9x+8

Found 2 solutions by aditya2117, MathTherapy:
Answer by aditya2117(32)

y=2x+7 ...................................................... (1)
y=9x+8 ...................................................... (2)
From (1),
x= (y-7)/2
Substituting value of x in eqn (2) we get,
y= 9.(y-7)/2 + 8
y = (9y-63+18)/2
2y=9y-45
9y-2y=45
7y=45
y=45/7
x=(45/7-7)/2=45-49/14=-4/14=-2/7
Reqd. Soln. : x = -2/7
y = 45/7

Answer by MathTherapy(10552)


solve the system of equations by the substitution method to get an ordered pair
y=2x+7
y=9x+8

y = 2x + 7 ------ eq (i)
y = 9x + 8 ------ eq (ii)
2x + 7 = 9x + 8 ------ Substituting 2x + 7 for y in eq (ii)
Solve for x, then substitute that value for x in any of the 2 equations to determine the value of y.
You can do the check!!
If you need a complete and detailed solution, let me know!!
Send comments, “thank-yous,” and inquiries to “D” at MathMadEzy@aol.com.
Further help is available, online or in-person, for a fee, obviously.